OpenStudy (anonymous):

The weight of a piece of crystal in the form of a regular dodecahedron is 360 lbs., and each of its edges is 10 inches long. What is the weight of a similar piece of crystal whose edge edge measures 16 in.?

hartnn (hartnn):

for same density, weight is proportional to volume and volume is proportional to cube of side. hence, weight is proportional to cube of side W1/W2 = S1^3/S2^3 360/W2 = (10/16)^3 u get W2=1474.56

OpenStudy (anonymous):

thanks!

hartnn (hartnn):

welcome ^_^